OpenStudy (anonymous):

What is the solution to the system of equations? 4x+y=2 x-y=3

OpenStudy (solomonzelman):

add the equations to each other, this will help you eliminate one parameter (the y).

OpenStudy (anonymous):

how do you add them?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 4x \quad+y\quad=2 }\) \(\large\color{black}{ \displaystyle ^{\huge ^+} \quad x\quad-y\quad=3 }\) \(\LARGE \color{black}{ \displaystyle ^\text{_____________} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 4x+x=\quad ? }\) \(\large\color{black}{ \displaystyle y+(-y)=\quad ? }\) \(\large\color{black}{ \displaystyle 2+3=\quad ? }\)

OpenStudy (anonymous):

4x+x= 4?

OpenStudy (solomonzelman):

Don't ask why I ask, just answer if you can, please.

OpenStudy (solomonzelman):

4 + 1 = ?

OpenStudy (anonymous):

5

OpenStudy (solomonzelman):

Just as: 4 +1 =5 So is: 4x +1x =5x

OpenStudy (solomonzelman):

(note that when you say "x" you really mean "1x") So, 4x+x=?

OpenStudy (anonymous):

5x

OpenStudy (solomonzelman):

Yes, and 3+2=?

OpenStudy (anonymous):

5

OpenStudy (solomonzelman):

Very nice...

OpenStudy (solomonzelman):

y - y = ?

OpenStudy (solomonzelman):

1 - 1 = 0 or, (Anything) - (Anything) = 0

OpenStudy (anonymous):

y+(-y)=0

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{red}{4x} \quad\color{green}{+y}\quad=\color{blue}{2} }\) \(\large\color{black}{ \displaystyle ^{\huge ^+} \quad \color{red}{x}\quad\color{green}{-y}\quad=\color{blue}{3} }\) \(\LARGE \color{black}{ \displaystyle ^\text{_____________} }\) \(\large\color{black}{ \displaystyle \color{red}{5x} \quad\color{green}{+0}\quad=\color{blue}{5} }\) <─ the result of adding these 2 equations

OpenStudy (solomonzelman):

Does everything make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

To say: 5x +0 = 5 Is the same as saying: 5x = 5 (Because adding zero makes no difference) And now you are able to solve for x (by dividng both sides by 5).

OpenStudy (anonymous):

x=1

OpenStudy (solomonzelman):

Very good!

OpenStudy (solomonzelman):

x=1. You can substitute this solution for x into any of the original equations to solve for y.

OpenStudy (solomonzelman):

Lets use the 2nd equation: x - y = 3 YOu are given that x=1, so you can re-write theis equation as: 1 - y =3

OpenStudy (solomonzelman):

Does everything make sense so far ? Can you solve for y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1-y=3 y = -2

OpenStudy (solomonzelman):

yes, x=1 and y=-2 Perfect!

OpenStudy (solomonzelman):

You made it !

OpenStudy (anonymous):

on the question i need hep with it is giving me cordintes that i have to fill in would that be (5x,5)

OpenStudy (solomonzelman):

Your coordinates are in the form of (x,y)

OpenStudy (solomonzelman):

Plug in the x-solution for x, and the y-solution for y.

OpenStudy (anonymous):

(5,5)

OpenStudy (anonymous):

is that correct

OpenStudy (solomonzelman):